### v1

# Automted Workload Mapping with 3D Heterogeneous Electronic-Photonic Processing-in-Memory Platform HPiMAP: A Framework for Mapping Hybrid DNNs onto Heterogeneous 3D-Stacked Electronic-Photonic Processing-in-Memory Architectures HPiMAP: A Multi-Objective DNN Mapping Framework on Heterogeneous Electronic-Photonic Processing-in-Memory Architectures

Abstract
The future of artificial intelligence (AI) acceleration demands a paradigm shift beyond the limitations of purely electronic or photonic architectures. Photonic analog computing delivers unmatched speed and parallelism but struggles with data movement, robustness, and precision. Electronic processing-in-memory (PIM) enables energy-efficient computing by co-locating storage and computation but suffers from endurance and reconfiguration constraints, limiting it to static weight mapping. Neither approach alone achieves the balance needed for adaptive, efficient AI. To break this impasse, we study a hybrid electronic-photonic-PIM computing architecture and introduce H3PIMAP, a heterogeneity-aware mapping framework that seamlessly orchestrates workloads across electronic and optical tiers. By optimizing workload partitioning through a two-stage multi-objective exploration method, H3PIMAP harnesses light speed for high-throughput operations and PIM efficiency for memory-bound tasks. System-level evaluations on language and vision models show H3PIMAP achieves a 2.74 energy efficiency improvement and a 3.47 latency reduction compared to homogeneous architectures and naïve mapping strategies. This proposed framework lays the foundation for hybrid AI accelerators, bridging the gap between electronic and photonic computation for next-generation efficiency and scalability.
I Introduction
Artificial Intelligence (AI), powered by deep neural networks (DNNs), has become integral to a wide array of application domains, e.g., computer vision, natural language processing (NLP), medical diagnostics, and time-series data analytics. These increasingly sophisticated workloads require low latency, high throughput, energy efficiency, and flexibility, objectives that traditional CPUs and GPUs often struggle to meet. To address issues in the notorious "memory wall" and growing computational inefficiency, researchers have explored various domain-specific accelerators that depart from the conventional von Neumann paradigm.
Among many emerging solutions, Processing-in-Memory (PIM) and photonic accelerator systems stand out. PIM harnesses the inherent capability of performing computations directly within or near the memory fabric [1], thereby circumventing the costly data transfers across separate processing and memory units. While PIM can substantially reduce latency and energy consumption, different memory technologies (e.g., SRAM [2], FeFET [3], ReRAM [4], MRAM [5]) each impose unique trade-offs involving footprint, endurance, write latency, and power dissipation. No single technology uniformly excels in all these dimensions, so effective PIM designs require careful technology and architectural selections.
Photonic accelerators, on the other hand, exploit the massive parallelism and high bandwidth of optical signals to process matrix-vector multiplications at unprecedented speeds. By encoding information in light, photonic computing can execute operations such as convolutions or fully connected layers with high throughput. Nonetheless, challenges like thermal crosstalk, process variations, limited data precision, and the overheads of optoelectronic conversion constrain the practical deployment of purely photonic systems. Developing robust calibration and error-compensation strategies remains a critical step toward unlocking the full potential of photonic accelerators.
Another motivation for embracing heterogeneous accelerators stems from the intrinsic heterogeneity in modern DNN workloads, for example, different layer sizes, synaptic sensitivity, input operand properties (dynamic or static), etc. These heterogeneous characteristics influence the suitability and efficiency of individual workload components (e.g., layers/channels) for mapping onto specific processing platforms. For example, mapping dynamic tensor products (attention layers in Transformers) to non-volatile PIM hardware [6] will cause significant endurance concerns. Fully leveraging the potential of heterogeneous accelerator platforms, therefore, demands an automated heterogeneity-aware workload mapping framework toward efficiency-optimal cooperation across hybrid AI accelerators.
In this work, we propose, for the first time, a heterogeneity-aware multi-objective mapping framework to intelligently distribute hybrid DNN workloads across heterogeneous electronic-photonic AI accelerators. Our two-stage mapping flow integrates state-of-the-art architecture and interconnect simulation to explore Pareto-optimal mapping solutions in energy-latency space. Then, our sensitivity-aware workload remapping stage rapidly boosts the mapping’s robustness against various hardware non-ideality to meet the accuracy target. Our key contributions are as follows:
Electronic-Photonic-PIM Accelerator Design: We introduce a novel hybrid accelerator that combines PIM hardware with photonics and achieves superior inference power, speed, and accuracy over homogeneous systems.
Electronic-Photonic-PIM Accelerator Evaluation Infrastructure: We create the first system modeling and optimization infrastructure integrating SoTA simulators for PIM, photonic accelerators, and network-on-chips, enabling automated design space exploration.
Heterogeneity-aware Mapping Framework: We introduce a two-stage multi-objective mapping framework (H3PIMAP) that considers workload variability and hardware heterogeneity, ensuring rapid mapping space exploration for optimal resource utilization, system performance, and inference accuracy.
Comprehensive Performance Evaluation: H3PIMAP shows 2.74 and 3.47 lower energy and latency over homogeneous systems and naive mapping strategy via evaluation over various NLP and Vision models/datasets.
II Related Work
We discuss relevant work that has been done within the scope of PIM and photonic DNN accelerator architectures. Specifically, we focus on homogeneous architectures solely based on either SRAM, ReRAM or photonic devices, reviewing their advantages and limitations. Table I compares the characteristics of SRAM, ReRAM, and photonic devices. Similar to our baseline ISAAC [7] and TeMPO [8] architectures, we discuss homogeneous PIM and photonic architectures and review prior work done on heterogeneous systems for DNN accelerators.
II-A PIM DNN Accelerators
PIM-based architectures offer an efficient solution for DNN inference by performing computations within the memory, thereby reducing memory traffic and data movement. DNN computations primarily involve matrix multiplications (MM), which are efficiently handled by the inherent parallelism in PIM crossbar arrays where the activations are applied directly to the word-lines and summed along the bit-lines of crossbars [7]. Due to the wide range of devices supporting PIM technology, significant differences in the performance, specifically energy efficiency and latency overhead, are observed. Each device also has its shortcomings. For instance, SRAM consumes more power [9], whereas ReRAM suffers from higher latency and increased device noise [10]. To address these trade-offs, our approach leverages the strengths of both memories to achieve optimal performance.
II-B Photonic Tensor Cores and Modeling Challenges
Photonic tensor cores (PTCs) are emerging hardware platforms for ultra-fast matrix multiplication using light [11, 12, 13]. The diverse properties of PTC designs result in variations in circuit topology, devices, and operational principles, making accurate hardware modeling challenging. For example, based on expressivity, universal PTCs can map arbitrary matrix multiplication [11, 12], while subspace PTCs support only a subset of static linear transformation [14]. Based on the numerical range of input operands, full-range PTCs handle arbitrary values in one-shot computation, while subspace coherent PTCs require multiple computations to obtain the same results. Reconfiguration speed is another key factor to determine its supported dataflow and operations. Certain PTCs require thermal tuning to reconfigure the weights, leading to delays of and limiting them to weight-stationary dataflows, which are unsuitable for dynamic self-attention. Dynamic PTCs [15] use high-speed modulators for real-time matrix switching, enabling dynamic tensor products and output-stationary dataflows. For this work, we focus on using dynamic PTCs as a complement to weight-static electronic PIM.
II-C Prior Heterogeneous Accelerators
Several works proposed DNN accelerators that exploit heterogeneity in PIM devices and digital architectures (e.g., GPUs, TPUs) to optimize performance, efficiency, and accuracy.
Significant efforts have also been made to mitigate non-idealities in ReRAM-based PIM.
For example, some designs employ SRAM for most-significant-bit calculations and ReRAM for least-significant-bit calculations following weight quantization [16], improving accuracy at the cost of higher power consumption and increased area.
Meanwhile, research such as HyDe [17] proposes to optimize the design with a single-objective (e.g., latency or accuracy) by identifying the optimal layer-to-device mapping, though it struggles in non-convex search regions.
Other works [18] have proposed a heterogeneous architecture that utilizes PIM for weight-stationary operations and TPUs for dynamic matrix multiplication in transformer models.
To the best of our knowledge, no existing research integrates photonics, which offers superior overall performance compared to TPUs, with PIM devices for DNN acceleration.
These gaps in research motivate us to introduce H3PIMAP, a mapping framework for efficient partitioning and mapping of DNN workloads to heterogeneous electronic-photonic-PIM accelerators.
III Proposed H3PIMAP Framework on Electronic-Photonic-PIM Architecture
We first illustrate the heterogeneous architecture with feature and noise analysis and then introduce our proposed mapping framework H3PIMAP for efficient, robust hybrid DNN workload mapping.
III-A Architecture Overview and Analysis
As a case study, we consider a multi-tier 3D-stacked accelerator architecture with ReRAM PIM tier (R), SRAM PIM tier (S), integrated photonics computing tier (P), and global buffer (M) tier, shown in Fig. 1. Each tier contains a multi-tile accelerator with one type of dedicated technology (R/S/P). Tiers are interconnected via silicon through vias (TSVs) for high-bandwidth data movement. The hierarchical on-chip global buffer for weights (needed by photonics) and activations (needed by all three computing tiers) is located at the memory tier (M).
As an important motivation, we answer a critical question: why do we need heterogeneous architectures that integrate three technologies (R/S/P)?
Feature Analysis of 3 Technologies – Key features of accelerators based on three technologies are summarized in Table I, highlighting key trade-offs that motivate the need for a heterogeneous architecture with a heterogeneity-aware mapping framework.
| Property | SRAM 22nm | ReRAM 32nm | Photonics |
| Resolution | 1-bit 8 cells=8-bit | 2-bit 4 cells = 8-bit | 46-bit |
| Tile Size | 256 crossbars, 128128 | 64 crossbars, 128128 | 2 cores, 1414 |
| ADC/tile | 256 SARADC 7-Bit | 64 SARADC 8-Bit | 392 SARADC 8-Bit |
| Cell Area | |||
| Arch Size | 100 Tiles | 100 Tiles | 2 Tiles |
| Program latency | 1 ns | 100 ns | 100 ps |
| Static power | Medium | Low | High |
| Clock | 100 MHz | 100 MHz | 3 GHz |
➊ Speed: Photonic accelerator exhibits unparalleled computational speed even with a few small-size tensor cores, achieving 110 TOPS performance. This makes them highly suitable for compute-bound workloads with high arithmetic intensity, e.g., convolution and self-attentions. In contrast, ReRAM-based PIM architectures typically show low operation speed due to their high latency and multiplexed ADC sampling mode, but excel in energy efficiency for memory-bound operations, i.e., large matrix multiplication operations, making them an ideal choice for efficiency-prioritized tasks. SRAM PIM accelerators fall between the other two technologies, offering balanced performance for moderately sized workloads.
➋ Weight-dynamic vs. Weight-static Operation Support: The inherent characteristics of these technologies dictate their suitability for different workload types. ReRAM and SRAM PIM architectures are well-suited for weight-static operations, where weights are mapped to dedicated PIM PEs to minimize frequent weight updates due to its weight-stationary dataflow and ReRAM endurance limits. On the other hand, dynamic operations, such as self-attention in Transformers, require frequent updates on both operands (e.g., Q and K tensors). are better handled by dynamic photonic PEs.
➌ Bit Resolution and Noise Robustness: SRAM and ReRAM PIM accelerators provide computation for 8-bit integer multiplication due to their high cell density and precise digital mechanisms. However, the 2-bit ReRAM cells are susceptible to thermal noises with conductance drift [10].
Both PIM architectures generate significant heat during operations, which can degrade the overall robustness of the PEs. This heat impacts not only the current tier where the computation is taking place but can also propagate to adjacent tiers in 3D-stacked architectures, compromising accuracy due to slow thermal dissipation. Photonic tensor cores, while capable of one-shot light-speed matrix multiplications, are typically limited to lower bit resolutions (e.g., <6-bit operands). Additionally, they are susceptible to hardware noise, which can affect computational accuracy.
III-B Dataflow and Interconnect Design
To match the distinct properties and operational demands of both PIM and photonic accelerators, we customize a dataflow to partition matrix multiplication in each DNN layer among three computing tiers (R/S/P). As each PIM tile communicates solely with the global buffer via TSVs and routers, data transfer remains one-dimensional. This approach eliminates inter-tile communication overhead. To further optimize communication efficiency, TSV connections to the PIM tiers are positioned midway between PIM tiles. This placement effectively halves the average communication distance relative to a 2D network-on-chip. Additionally, we add a dedicated TSV link to the photonic tier, accommodating its substantial bandwidth requirements. Our following introduced mapping algorithm and system evaluations are based on the above dataflow and interconnect designs.
III-C Non-ideal Hardware Noise Modeling
As our architecture contains analog computing hardware (ReRAM and photonics), their hardware non-ideality, e.g., noise variation, will impact the computing fidelity and model inference accuracy. Their unique noise properties imply that a noise-aware mapping strategy can boost inference accuracy. To understand the hardware robustness, we utilize standard models to capture device-level noise variations across different tiers.
PIM: We consider SRAM to be robust to thermal noises [9] for the scope of this work due to its high thermal tolerance and digital computing mechanism. However, ReRAM is thermally sensitive. Major sources of noise in ReRAM [10] are thermal and shot noise, formulated as
(1) where , , , , and denote the conductance, terminal voltage, Boltzmann constant, temperature, and frequency, respectively.
Photonics: We adopt the TeMPO architecture [8] and its real-device noise measurements as our noise model. Following its original settings, the photonic noise is treated as a relative random Gaussian perturbation injected into both the input tensors of matrix multiplications, formally modeled as , where . According to TeMPO, real measurements under typical operating conditions show that is most representative, and thus we adopt as photonic tier noise.
III-D Mapping Problem Formulation
Given an -layer DNN with sequential layer execution, our goal is to determine an optimal matrix row-to-tier mapping individually for each layer on our heterogeneous architecture. Each weight matrix row is assigned to one of the computing tiers. We use to represent the percentage of matrix rows mapped to tier . The objective of the mapping process is to minimize the overall model inference energy () and latency () under several specific constraints.
This can be formulated as the following multi-objective optimization (MOO) problem:
The final configuration specifies how each layer’s weight rows are assigned across tiers. The total inference latency equals the sum of individual layer latencies, with each layer bottlenecked by the slowest tier. The overall energy consumption is computed by adding the energy usage of all layers. Here, denotes the total memory capacity (or tile size) of tier . To avoid storage overflow, no tier may be assigned more weight rows than it can accommodate. As discussed in ➋, any operation that a given tier cannot support will not be mapped to that tier. Because device noise and limited precision can degrade performance, we require to remain within discrepancy below the original accuracy .
Search Space Analysis: The complexity of the problem can be illustrated with the following example. For an -layer DNN with neurons (weight rows) per layer mapped onto tiers, there are total possible mappings. For example, a popular language model Pythia-70M has an average of 2048 neurons per layer and 6 layers in total. With 3 tiers to map, the total search space is 312288. Besides the vast search space, the non-trivial cost of evaluating the accuracy and hardware efficiency of each solution cast significant challenges to the MOO task. To efficiently explore the design space of this constrained multi-objective optimization problem, we propose H3PIMAP, a two-stage search and optimization procedure as shown in Fig. 2. We decouple the accuracy and efficiency optimization in two stages and prune the search space with sensitivity-aware heuristics. The first stage rapidly explores the energy-latency space to obtain Pareto-optimal mapping candidates. Then, the second stage performs accuracy-driven row remapping to strategically calibrate the row assignment to satisfy the accuracy constraints.
III-D1 Stage 1: Latency-Energy Pareto Optimization (PO)
In this stage, we perform a multi-objective search that focuses on exploring the Pareto front of energy–latency trade-offs. Since only the number of rows mapped to a tier impacts the inference latency and energy, not specific row indices, we are allowed to significantly reduce the exponential mapping space. The pruned search space now reduces from to . For Pythia-70M, the space is reduced from to 8.61037.
We employ a tailored evolutionary algorithm NSGA-II [19] to solve stage 1, which queries the tier-specific architecture simulators to get the fitness for a particular mapping. The algorithm of stage 1 is detailed in Alg. 1. The simplified energy-latency Pareto optimization problem is
Once we obtain the Pareto-front populations with high efficiency and speed, we examine whether the best-accuracy solution among them meets the accuracy constraint. If such a solution exists, the best-accuracy one will be the final solution. If not, we will pass the best-accuracy one to the second stage to adjust the mapping until the accuracy target is met. Due to device-specific bit precision and noise, we strategically use row-wise sensitivity to guide the assignment. Specifically, we compute the sensitivity of each row using a second-order Taylor expansion with Gaussian perturbations:
where is the loss function, is the gradient of the loss with respect to one row in layer , the hessian matrix approximated from its diagonal entirety , and represents perturbations in row of the weight matrix. A sorted tier from best to worst model performance can be acquired by measuring the precision of each tier on the same workload. Next, using the row-wise sensitivity information, we perform a sorted assignment between the rows and tiers, i.e., mapping the most sensitive rows to the most accurate tiers.
Finally, we evaluate the configuration to ensure the overall accuracy meets the required threshold. If so, we have a valid mapping configuration that satisfies both performance requirements and accuracy. Otherwise, we choose the mapping with the highest model performance () and proceed to the second stage, monotonic optimization, to further improve accuracy.
III-D2 Stage 2: Acc-Driven Row-Remap (RR)
In the accuracy optimization stage, we iteratively reassign weight rows from the most noise-prone tier to the tier offering higher accuracy to ensure the process converges to the required accuracy. Although the initial weight mapping (acquired in Stage 1) provides a per-layer distribution, we convert it into a global distribution across all tiers for the entire model. Throughout the reassignment process, we continuously verify that no tier’s memory usage exceeds its capacity. The algorithm finalizes and returns the current assignment once the best tier’s memory is fully occupied or the accuracy requirement is satisfied.
IV Evaluation Results
IV-A Evaluation Setup
Our experiments use the Pythia-70M model [23] trained on the TinyStories dataset [24] as our baseline for studies and analysis. In addition, we evaluate the MobileViT-S model [25] on two distinct datasets: a military asset [26] and a medical chest X-ray dataset [27] to prove our framework’s versatility. All models and datasets are shown in Table II and III.
| Datasets | Dataset Size | Training Set | Test Set | Benchmark |
| TinyStories | 1.92GB | 2.12M rows | 22K rows | 1.1017 (PPL) |
| Military Assets | 4.19 GB | 21978 samples | 1396 samples | 0.8972 (Acc) |
| Chest X-Ray Images | 1.24 GB | 5216 samples | 624 samples | 0.9533 (Acc) |
| Models | Params |
| MobileViT-S | 5.6 M | 69 | 37 | 32 | 9 | 18 | ||||||||||
| Pythia | 70 M | 24 | 24 | 0 | 6 | 12 |
The models we evaluate retain their original architectures, modified only by including learned step quantization [28]. We first train all models from scratch in an 8-8-8-bit (input–weight–output) configuration, then fine-tune a 6-6-8-bit variant from the 8-bit checkpoint. This fine-tuning step helps maintain a smooth distribution when moving to lower bit precision, accommodating tiers with constrained precision. Note that this fine-tuned 6-6-8-bit model is only used in the RR stage. All experiments are performed on a Linux OS server with NVIDIA A6000 GPUs and AMD EPYC 7763 64-core processors.
IV-B Results
IV-B1 Network-on-Chip Interconnect Cost
Using BookSim NoC simulator [22], we simulate the performance of 2.5D and 3D topologies with inputs of size [8, 3, 32, 32] between two Conv2D layers, shown in Fig. 3. The experiment shows an improvement of 40% and 41% in latency and energy cost, respectively.
IV-B2 Pareto Optimization (PO)
We begin our optimization process by initializing the population and then allowing the NSGA-II [19] algorithm to search for Pareto-optimal solutions of mapping Pythia-70M. Figure 4 illustrates how energy and latency are jointly optimized throughout this evolutionary search. Figure 5 ➊ shows the layer-wise workload distribution among three tiers (Left) and the row-wise mapping results of each layer (Right) after PO.
IV-B3 Accuracy-Driven Row-Remapping (RR)
| Strategy | Latency (ms) | Energy (mJ) |
Latency (ms) | Energy (mJ) |
Latency (ms) | Energy (mJ) |
| 100% SRAM | 10.21 | 13.79 | 1.1017 | 291.92 | 4.70 | 0.953 | 291.92 | 4.70 | 0.897 | ||||||
| 100% ReRAM | 14.73 | 13.44 | 1.1128 | 583.54 | 3.97 | 0.949 | 583.54 | 3.97 | 0.888 | ||||||
| 100% TeMPO | 0.91 | 8.92 | 2.2272 | 2.17 | 15.71 | 0.711 | 2.174 | 15.71 | 0.507 | ||||||
[…]
Once we obtain the set of Pareto-optimal solutions, there are inevitably multiple candidates that balance our objectives in different ways. To narrow this down, we explicitly evaluate the accuracy for each of these candidates and select the one that yields the best final accuracy. As a preliminary step, we calculated each row’s sensitivity using a Taylor expansion approach using Eq. (4). Figure 5 right side shows the row-wise sensitivity distribution of each layer sorted from the most sensitive layer to the least one. Leveraging these sensitivity insights, we then gradually shift a portion of the photonic chip’s workload towards SRAM. Since SRAMs are not affected by low-precision and noise-related issues, this rebalancing can help retain the model’s accuracy. As shown in Fig. 6, the PPL steadily improves as the optimization progresses. We impose a 0.1 tolerance relative to the noise-free PPL for RR. Comparing the final workload assignment after RR (Fig. 5 ➋) with the one in PO (Fig. 5 ➊), we observe a noticeable shift in workload from the Photonic tier to SRAM.
| Strategy | Latency (ms) | Energy (mJ) |
Perplexity | LEP Score | ||
| 100% SRAM | 10.21 | 13.79 | 8-8-8 | 1.1017 | 0.5580 | ||
| 100% ReRAM | 14.73 | 13.44 | 8-8-8 | 1.1128 | 0.6428 | ||
| 100% TeMPO | 0.91 | 8.92 | 6-6-8 | 2.2272 | 0.3333 | ||
| Equal Distribution | 4.90 | 12.02 | Mixed | 1.1861 | 0.3339 | ||
[…]
IV-B4 Heterogeneous vs. Homogeneous
To validate our approach, we compare three methods: (1) our proposed framework, (2) an intuitive “equal workload” partition, and (3) a homogeneous workload assignment. Results are shown in Table V. From the table, PO achieves superior efficiency: it reduces latency by 3.66 and energy consumption by 1.22 compared to the equal-workload method, and by 6.42 on latency and 1.22 on energy on average compared to the homogeneous assignment. However, under the PO workload assignment, the model’s performance does not meet the 0.1-difference constraint; after RR, it improves substantially while still maintaining a great latency and energy profile. To facilitate a comprehensive comparison, we introduce a Latency–Energy–Performance (LEP) score, which normalizes and averages three metrics such that lower scores indicate better overall efficiency. As shown in Table V, PO achieves the lowest LEP score; following RR, performance increases more while retaining a second low LEP value. Figure 7 shows each layer’s latency and energy.
IV-B5 Main Results
Lastly, we compare the result of H3PIMAP mapping Pythia-70M (trained on TinyStories) and MobileViT-S (trained on Chest X-Ray and Military Assets) to the proposed 3D heterogeneous architecture with homogeneous mapping, shown in Table IV. Homogeneous mapping shows excessive latency overhead in PIM tiers and unacceptable accuracy degradation in the photonic tier. H3PIMAP mapping results show the solution where the latency and energy are lowered by 3.47 and 2.74, respectively, while maintaining a comparable model accuracy.
V Conclusion
This work presents H3PIMAP, a heterogeneity-aware mapping framework that optimally distributes hybrid DNN workloads across a novel Electronic-Photonic-PIM heterogeneous AI computing architecture. By integrating multi-objective optimization with accuracy-driven remapping, H3PIMAP achieves a Pareto-optimal balance of latency and energy efficiency while ensuring robust model inference against hardware non-ideal variations. On vision and language benchmarks, H3PIMAP achieves a 3.47 reduction in latency and a 2.74 improvement in energy efficiency compared to standard mapping on homogeneous systems. These results highlight the transformative potential of heterogeneous AI hardware, where the synergistic integration of electronic PIM and photonics, coupled with efficient software workload mapping, unlocks new frontiers in next-generation AI acceleration.